Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Mr. Yutaro’s class contains 5 boys and 11 girls

Expert replies
by BTGmoderatorDC » Wed Dec 27, 2017 12:50 am
Mr. Yutaro’s class contains 5 boys and 11 girls. If two students are chosen, one at a time, from the class, what is the probability that a boy and a girl are chosen?

A. 9/48
B. 9/20
C. 11/24
D. 27/48
E. 13/16

Can some experts show me a simple solution in this problem?

OA C
Join the discussion
Source: — Problem Solving |

by elias.latour.apex » Wed Dec 27, 2017 4:58 am
There are several ways to solve this problem. One way is to calculate the chance that a boy is chosen first and then a girl is chosen second. This probability will be (5/16)(11/15) but this is only half the probability because we must also consider the chance that a girl is chosen first and then a boy. That probability is (11/16)(5/15). The odds are the same; only the numerators have changed places. Thus, we can simply calculate the probability as 2(11/16)(5/15).

This reduces to (11/8)(1/3) so the chances are 11/24, which is answer choice C.
Elias Latour
Verbal Specialist @ ApexGMAT
blog.apexgmat.com
+1 (646) 736-7622
Join the discussion

by Brent@GMATPrepNow » Thu Dec 28, 2017 3:32 pm
lheiannie07 wrote:Mr. Yutaro’s class contains 5 boys and 11 girls. If two students are chosen, one at a time, from the class, what is the probability that a boy and a girl are chosen?

A. 9/48
B. 9/20
C. 11/24
D. 27/48
E. 13/16
P(1 boy and 1 girl chosen) = P(1st selection is boy AND 2nd selection is girl OR 1st selection is girl AND 2nd selection is boy)
= P(1st selection is boy AND 2nd selection is girl) + P(1st selection is girl AND 2nd selection is boy)
= P(1st selection is boy) x P(2nd selection is girl) + P(1st selection is girl) x P(2nd selection is boy)
= 5/16 x 11/15 + 11/16 x 5/15
= 55/240 + 55/240
= 110/240
= 11/24
= C

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by GMATGuruNY » Thu Dec 28, 2017 10:24 pm
lheiannie07 wrote:Mr. Yutaro’s class contains 5 boys and 11 girls. If two students are chosen, one at a time, from the class, what is the probability that a boy and a girl are chosen?

A. 9/48
B. 9/20
C. 11/24
D. 27/48
E. 13/16
P = (good outcomes)/(all possible outcomes).

All possible outcomes:
From 16 students, the number of ways to choose 2 = 16C2 = (16*15)/(2*1) = 120.

Good outcomes:
A good outcome consists of 1 boy and 1 girl.
Of the 16 students, 5 are boys, 16 are girls.
Thus:
Number of options for the boy = 5.
Number of options for the girl = 11.
To combine these options, we multiply:
5*11 = 55.

Resulting probability:
(good outcomes)/(all possible outcomes) = 55/120 = 11/24.

The correct answer is C.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by Scott@TargetTestPrep » Tue Sep 17, 2019 8:41 am
BTGmoderatorDC wrote:Mr. Yutaro’s class contains 5 boys and 11 girls. If two students are chosen, one at a time, from the class, what is the probability that a boy and a girl are chosen?

A. 9/48
B. 9/20
C. 11/24
D. 27/48
E. 13/16

Can some experts show me a simple solution in this problem?

OA C
The probability that a girl and a boy, in that order, are chosen is 11/16 x 5/15 = 11/16 x 1/3 = 11/48. The probability that a boy and a girl, in that order, are chosen is also 11/48. Therefore, the probability that a boy and a girl are chosen is 11/48 + 11/48 = 22/48 = 11/24.

Answer: C

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion